Site Conditions:

  • Smoke cannot be observed on the site
  • Odor can be detected throughout the site, more noticeable near the entrance

Activities:

  • Over the weekend, the majority of the northern area, Section B, was seeded; expected to begin spreading hay in the area on Tuesday
  • The southern area, Section A, has been completely seeded and covered with hay; grass can be seen continuously along the slope
  • Plan to place aggregate along the edges of the site today as BMP
  • The offsite borrow pit will begin closing operations today; the area will have hay and seed put down over the next week
  • EPA plans to continue the air monitoring until March 24